OpenStudy (anonymous):

lim as x approches 0 (sin^2X)/(X)

OpenStudy (zarkon):

0

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{\sin^2(x)}{x}=\lim_{x\to 0}\sin(x)\frac{\sin(x)}{x}=0\cdot1=0\]